how many sides are in a polygon if its interior angles sum up to 1,800 degrees, n=?

Answer by Alan3354(31487)

how many sides are in a polygon if its interior angles sum up to 1,800 degrees, n=?
------------
Sum of the angles = 180*(n-2)
180*(n-2) = 1800
n-2 = 10
n = 12